# Swarm-based gradient descent meets simulated annealing

### v1

@@ 2.2. Large-time convergence— from mean-field to global minimum
Observe that even if ${\mathcal{N}}_{\epsilon}\cap\mathrm{supp}_{\mathbf{x}}\mu_{0}(\cdot,m)=\emptyset$, then the presence of the Brownian motion implies that there exists a time, $t\in(0,1]$, such that ${\mathcal{N}}_{\epsilon}\cap\mathrm{supp}_{\mathbf{x}}\mu_{t}(\cdot,m)\neq\emptyset$, and consequently, (2.9) always yields a finite number threshold, $t_{\epsilon}$. However, we do not claim that the ‘convergence rate’ worked out below, $t_{\epsilon}\lesssim\epsilon^{-1}$, is optimal: our numerical findings reported in Section 6 indicate the possibility of a much better convergence rate in the sense that $t_{\epsilon}=\log(1/\epsilon)$ is already sufficient to ensure (2.10).
@@
Assume that Assumptions 2.1, 2.2 hold. Let $\mu^{N}_{t}(\mathbf{x},m)$ be the empirical distribution associated with an ensemble of solutions of the swarm-based optimization with adjusted annealing rate  (1.2), $(\mathbf{x}^{j}_{t},m^{j}_{t})$, subject to the initial data drawn i.i.d. from $\mu_{t=0}$. Fix an arbitrarily $\epsilon>0$. Then, for a large enough crowd spelled out in (2.6), depending on time, $t_{\epsilon/2}$ defined in (2.9),
@@ 3. From mean-field to macroscopic description
We shall briefly comment on the macroscopic description of the swarm-based optimization with adjusted annealing rate, furnished in terms of the firsts two moments — the density, $\rho:{\mathbb{R}}_{+}\times{\mathbb{R}}^{d}\mapsto{\mathbb{R}}_{+}$, and momentum, $\rho{\sf v}:{\mathbb{R}}_{+}\times{\mathbb{R}}^{d}\mapsto{\mathbb{R}}$, (observe that time-dependence in macroscopic quantities is denoted as usual $\rho=\rho(t,\mathbf{x}),{\sf v}={\sf v}(t,\mathbf{x})$ etc.)
$\rho(t,\mathbf{x}):=\int\mu_{t}(\mathbf{x},m)\textnormal{d}m,\qquad\rho{\sf v}(t,\mathbf{x}):=\int m\mu_{t}(\mathbf{x},m)\textnormal{d}m.$
@@
(3.1) $\left\{\begin{split}\rho_{t}+\nabla_{\mathbf{x}}\cdot(\rho\nabla F)&=\Delta_{\mathbf{x}}\int\sigma(m)\mu_{t}(\mathbf{x},m)\textnormal{d}m,\\ (\rho{\sf v})_{t}+\nabla_{\mathbf{x}}\cdot(\rho{\sf v}\nabla F)&=\big{(}\overline{F}(t)-F(\mathbf{x})\big{)}\rho{\sf v}+\Delta_{\mathbf{x}}\int m\sigma(m)\mu_{t}(\mathbf{x},m)\textnormal{d}m.\end{split}\right.$
We normalize the total mass and momentum, $\int\rho(t,\mathbf{x})\textnormal{d}\mathbf{x}\equiv 1$ and $\int\rho{\sf v}(t,\mathbf{x})\textnormal{d}\mathbf{x}\equiv 1$, which leaves us with a provisional minimum, $\overline{F}(t)=\overline{\mathscr{F}}^{\mu}_{t}$, given by
$\overline{F}(t)=\int F(\mathbf{x})\rho{\sf v}(t,\mathbf{x})\textnormal{d}\mathbf{x}.$
To unlock a closed form of the diffusion on the right, we assume the simplest closure based on a mono-kinetic pseudo-Maxwellian , $\mu_{t}(\mathbf{x},m)=\rho(t,\mathbf{x})\delta(m-{\sf v}(t,\mathbf{x}))$. This yields
(3.2a) $\rho_{t}+\nabla_{\mathbf{x}}\cdot(\rho\nabla F)$ $=\Delta_{\mathbf{x}}(\sigma(t,\mathbf{x})\rho(t,\mathbf{x})),\qquad\sigma(t,\mathbf{x}):=\sigma({\sf v}(t,\mathbf{x}))$
(3.2b) $(\rho{\sf v})_{t}+\nabla_{\mathbf{x}}\cdot(\rho{\sf v}\nabla F)$ $=\big{(}\overline{F}(t)-F(\mathbf{x})\big{)}\rho{\sf v}+\Delta_{\mathbf{x}}\big{(}\sigma(t,\mathbf{x})\rho{\sf v}(t,\mathbf{x})\big{)}.$
The momentum (3.2b) can be converted into a drift-diffusion equation for the velocity ${\sf v}$,
(3.3a) $\rho_{t}+\nabla_{\mathbf{x}}\cdot(\rho\nabla F)$ $=\Delta_{\mathbf{x}}(\sigma(t,\mathbf{x})\rho(t,\mathbf{x})),$
(3.3b) ${\sf v}_{t}+\nabla F\cdot\nabla_{\mathbf{x}}{\sf v}$ $=\big{(}\overline{F}(t)-F(\mathbf{x})\big{)}{\sf v}+\frac{1}{\rho}\Delta_{\mathbf{x}}\big{(}\sigma(t,\mathbf{x})\rho{\sf v}(t,\mathbf{x})\big{)},\quad\mathbf{x}\in\textnormal{supp}\{\rho(t,\cdot)\}.$
@@
with expected long-time behavior $\rho(t,\mathbf{x})\stackrel{{\scriptstyle t\rightarrow\infty}}{{\longrightarrow}}\delta(\mathbf{x}-\mathbf{x}_{*})$ and ${\sf v}(t,\mathbf{x})\stackrel{{\scriptstyle t\rightarrow\infty}}{{\longrightarrow}}{\mathds{1}}(\mathbf{x}_{*})$. Observe that $\sigma(t,\mathbf{x})$ should vanish as $\mathbf{x}\rightarrow\mathbf{x}_{*}$.
@@ 4.1. Swarming with no communication
What is the effect of Brownian motion in our swarm-based optimization with adjusted annealing rate? If we turn off the amplitude, $\sigma(m)\equiv 0$, then in (1.2) is reduced to a deterministic system of non-interacting agents (and again we switch notations of time-dependence in deterministic quantities, $\mathbf{x}^{j}(t),m^{j}(t),\overline{F}^{N}(t)$ etc.)
@@
The essential role of communication was already highlighted in the context of SBGD method, [21]: once we decouple the step-size and mass, setting a uniform step-size in (1.1), that is, $\alpha(m)\equiv 1$, then the success rate of the resulting non-communicating swarm-based method decreases dramatically. The system (4.1) provides yet another realization of the essential role of communication. We note that since all particles are “descending” it is also straightforward to show that the collective mean objective $\overline{F}^{N}(t)$ is always decreasing: $\frac{\textnormal{d}}{\textnormal{d}t}\overline{F}^{N}(t)\leqslant 0$. See Section A.1 for the proof. However, even though $\overline{F}^{N}(t)$ continues to decay in time, it does not guarantee the desired convergence to the global minimum, $\liminf_{t\rightarrow\infty}\overline{F}^{N}(t)=\min_{\mathbf{x}}F(\mathbf{x})$. This is due to the passive nature of the gradient descent of the agents. In particular, it is likely that all agents will become trapped at local minima or saddle points, especially when the number of particles is not sufficiently large. Specifically, we have the following.
@@ 5.1. Convergence to the mean-field limit
Theorem 2.1 states that the empirical distribution $\mu^{N}_{t}$ that assembles all samples that follow the coupled SDE system (1.2), is close to the limiting distribution $\mu$ that solves the PDE (2.4a).
There are a few different but related approaches for justifying mean-field limits, including the coupling method, the energy/entropy estimates, the utilization of BBGKY hierarchies and running the tightness argument to confine a sequence of measures for the convergence; see references [22, 27]. Here we employ the coupling method, introduced in [27]. The coupling method places a stronger requirement on the field, but the machinery is rather simple to use, and it has the advantage of providing precise convergence rate in terms of the number of particles, $N$.
@@
(5.2) $\widetilde{\mu}^{N}_{t}=\frac{1}{N}\sum_{i}\delta_{\widetilde{\mathbf{x}}^{j}_{t}}(\mathbf{x})\otimes\delta_{\widetilde{m}^{j}_{t}}(m)\,,\quad\text{and}\quad\widetilde{F}^{N}_{t}=\frac{\sum_{j}\widetilde{m}^{j}_{t}F\left(\widetilde{\mathbf{x}}^{j}_{t}\right)}{\sum_{j}\widetilde{m}^{j}_{t}}.$
We note that this system has a clear similarity with the original SDE (1.2) but is passive, in the sense that the dynamics of $\widetilde{m}^{j}_{t}$ is driven by ¯$\overline{\mathscr{F}}^{\mu}$ instead of $\widetilde{F}^{N}_{t}$. While the dynamics of $\widetilde{\mathbf{x}}_{t}$ agrees with that of $\mathbf{x}_{t}$, the mass changes call for different rates. In particular, in (1.2), we compare $F(\mathbf{x}_{t})$ with $\overline{F}^{N}_{t}$ that is self-generated in (2.2), and here, $F(\widetilde{\mathbf{x}}_{t})$ is compared with ¯$\overline{\mathscr{F}}^{\mu}$ defined from the underlying PDE system (2.4a). As a consequence, this new $\{\widetilde{\mathbf{x}}^{j}_{0}=\mathbf{x}^{j}_{0}\,,\widetilde{m}^{j}_{0}=m^{j}_{0}\}$ system gets passively pushed forward by the underlying $\mu$.
@@
    $\lim_{N\to\infty}\mathbb{E}(W_{2}(\mu_{t},\widetilde{\mu}^{N}_{t}))=0.$
@@
Suppose Assumption 2.1 holds. Let $\{\widetilde{\mathbf{x}}^{j}_{t}\,,\widetilde{m}^{j}_{t}\}_{j=1}^{N}$ be the system defined in (5.1), and let $\{{\mathbf{x}}^{j}_{t}\,,{m}^{j}_{t}\}_{j=1}^{N}$ solve (1.2). Suppose the two systems have the same initial conditions, then
@@
    $\lim_{N\to\infty}\mathbb{E}(W_{2}(\mu^{N}_{t},\widetilde{\mu}^{N}_{t}))=0$
@@
    $\mathbb{E}(|\widetilde{F}^{N}_{t}-\overline{F}^{N}_{t}|)<\frac{C}{\sqrt{N}}\,.$
@@
With these two lemmas, Theorem 2.1 is a straightforward corollary, calling triangle inequality:
$\mathbb{E}\left(W_{2}(\mu_{t},\mu^{N}_{t})\right)\leqslant\mathbb{E}\left(W_{2}(\mu_{t},\widetilde{\mu}^{N}_{t})\right)+\mathbb{E}\left(W_{2}(\widetilde{\mu}^{N}_{t},\mu^{N}_{t})\right)\,.$
We omit the details here and move on to provide the proof of the two lemmas.
Proof for Lemma 5.1.
Calling [9],
@@
with $\theta=\max\{1/2,1/d\}$.
@@
The proof for Lemma 5.2 is the standard coupling method. This amounts to compare the original SDE (1.2) with the auxiliary SDE (5.1). In particular, we note that the $\mathbf{x}$ component of the two systems are almost identical:
@@
Proof for Lemma 5.2.
Noting that $\mu^{N}_{t}=\frac{1}{N}\sum_{i}\delta_{\mathbf{x}^{j}_{t}}(\mathbf{x})\otimes\delta_{m^{j}_{t}}(m)$ and $\widetilde{\mu}^{N}_{t}=\frac{1}{N}\sum_{i}\delta_{\widetilde{\mathbf{x}}^{j}_{t}}(\mathbf{x})\otimes\delta_{\widetilde{m}^{j}_{t}}(m)$, then recall the definition of the Wasserstein metric, we have:
@@
Similarly, for the mean objective, we have, according to the definition:

### v2

@@ 2.2. Large-time convergence— from mean-field to global minimum
Observe that even if ${\mathcal{N}}_{\epsilon}\cap\mathrm{supp}_{\mathbf{x}}\mu_{0}(\cdot,m)=\emptyset$, then the presence of the Brownian motion implies that there exists a time, $t\in(0,1]$, such that ${\mathcal{N}}_{\epsilon}\cap\mathrm{supp}_{\mathbf{x}}\mu_{t}(\cdot,m)\neq\emptyset$, and consequently, (2.9) always yields a finite number threshold, $t_{\epsilon}$. The combination of (2.9)-(2.10) together claims that with roughly $t_{\epsilon}\approx\frac{1}{\epsilon}$ time, the global optimal value can be found within $\epsilon$ accuracy, implying an $\mathcal{O}(1/\epsilon)$ convergence. However, the constant dependence heavily rely on the initial data, and we cannot claim the optimality of the ‘convergence rate’: our numerical findings reported in Section 6 indicate the possibility of a much better convergence rate in the sense that $t_{\epsilon}=\log(1/\epsilon)$ is already sufficient to ensure (2.10). See Figure 6.4.
@@
Assume that Assumptions 2.1, 2.2 hold [^4] Assumption 2.2 requires the uniqueness of the global minimum. This assumption can be replaced by finite many global minima $\{\mathbf{x}_{k}\}_{k=1}^{K}$ with $F(\mathbf{x}_{k})=F_{*}$ for all $k$. The details of the analysis need to be revised accordingly. In the proof (Section 5.2), we point out the specific location to revise to accommodate the situation where multiple global minima co-exist.. Let $\mu^{N}_{t}(\mathbf{x},m)$ be the empirical distribution associated with an ensemble of solutions of the swarm-based optimization with adjusted annealing rate  (SSA), $(\mathbf{x}^{j}_{t},m^{j}_{t})$, subject to the initial data drawn i.i.d. from $\mu_{t=0}$. Fix an arbitrary $\epsilon>0$. Then, for a large enough crowd spelled out in (2.6), depending on time, $t_{\epsilon/2}$ defined in (2.9),
@@ 3. From mean-field to macroscopic description
We shall briefly comment on the macroscopic description of the swarm-based optimization with adjusted annealing rate, furnished in terms of the first two moments — the density, $\rho:{\mathbb{R}}_{+}\times{\mathbb{R}}^{d}\mapsto{\mathbb{R}}_{+}$, and momentum, $\rho{\sf M}:{\mathbb{R}}_{+}\times{\mathbb{R}}^{d}\mapsto{\mathbb{R}}$, (observe that time-dependence is denoted as usual for the density $\rho=\rho(t,\mathbf{x})$, macroscopic mass, ${\sf M}={\sf M}(t,\mathbf{x})$ etc.)
$\rho(t,\mathbf{x}):=\int\mu_{t}(\mathbf{x},m)\textnormal{d}m,\qquad\rho{\sf M}(t,\mathbf{x}):=\int m\mu_{t}(\mathbf{x},m)\textnormal{d}m.$
@@
(3.1) $\left\{\begin{split}\rho_{t}-\nabla_{\mathbf{x}}\cdot(\rho\nabla F)&=\Delta_{\mathbf{x}}\int\sigma(m)\mu_{t}(\mathbf{x},m)\textnormal{d}m,\\ (\rho{\sf M})_{t}-\nabla_{\mathbf{x}}\cdot(\rho{\sf M}\nabla F)&=\big{(}\overline{F}(t)-F(\mathbf{x})\big{)}\rho{\sf M}+\Delta_{\mathbf{x}}\int m\sigma(m)\mu_{t}(\mathbf{x},m)\textnormal{d}m.\end{split}\right.$
We normalize the total mass and momentum, $\int\rho(t,\mathbf{x})\textnormal{d}\mathbf{x}\equiv 1$ and $\int\rho{\sf M}(t,\mathbf{x})\textnormal{d}\mathbf{x}\equiv 1$, which leaves us with a provisional minimum, $\overline{F}(t)=\overline{\mathscr{F}}^{\mu}_{t}$, given by
$\overline{F}(t)=\int F(\mathbf{x})\rho{\sf M}(t,\mathbf{x})\textnormal{d}\mathbf{x}.$
To unlock a closed form of the diffusion on the right, we assume the simplest closure based on a mono-kinetic pseudo-Maxwellian , $\mu_{t}(\mathbf{x},m)=\rho(t,\mathbf{x})\delta(m-{\sf M}(t,\mathbf{x}))$. This yields
(3.2a) $\rho_{t}-\nabla_{\mathbf{x}}\cdot(\rho\nabla F)$ $=\Delta_{\mathbf{x}}(\sigma(t,\mathbf{x})\rho(t,\mathbf{x})),\qquad\sigma(t,\mathbf{x}):=\sigma({\sf M}(t,\mathbf{x}))$
(3.2b) $(\rho{\sf M})_{t}-\nabla_{\mathbf{x}}\cdot(\rho{\sf M}\nabla F)$ $=\big{(}\overline{F}(t)-F(\mathbf{x})\big{)}\rho{\sf M}+\Delta_{\mathbf{x}}\big{(}\sigma(t,\mathbf{x})\rho{\sf M}(t,\mathbf{x})\big{)}.$
The momentum (3.2b) can be converted into a drift-diffusion equation for the velocity ${\sf M}$,
(3.3a) $\rho_{t}-\nabla_{\mathbf{x}}\cdot(\rho\nabla F)$ $=\Delta_{\mathbf{x}}(\sigma(t,\mathbf{x})\rho(t,\mathbf{x})),$
(3.3b) ${\sf M}_{t}-\nabla F\cdot\nabla_{\mathbf{x}}{\sf M}$ $=\big{(}\overline{F}(t)-F(\mathbf{x})\big{)}{\sf M}+\frac{1}{\rho}\Delta_{\mathbf{x}}\big{(}\sigma(t,\mathbf{x})\rho{\sf M}\big{)}-\frac{1}{\rho}\Delta_{\mathbf{x}}\big{(}\sigma(t,\mathbf{x})\rho\big{)}{\sf M},\quad\mathbf{x}\in\textnormal{supp}\{\rho(t,\cdot)\}.$
@@
with expected long-time behavior $\rho(t,\mathbf{x})\stackrel{{\scriptstyle t\rightarrow\infty}}{{\longrightarrow}}\delta(\mathbf{x}-\mathbf{x}_{*})$ and ${\sf M}(t,\mathbf{x})\stackrel{{\scriptstyle t\rightarrow\infty}}{{\longrightarrow}}{\mathds{1}}(\mathbf{x}_{*})$. Observe that $\sigma(t,\mathbf{x})$ should vanish as $\mathbf{x}\rightarrow\mathbf{x}_{*}$.
@@ 4.1. Swarming with no communication
What is the effect of Brownian motion in our swarm-based optimization with adjusted annealing rate? If we turn off the amplitude, $\sigma(m)\equiv 0$, then (SSA) is reduced to a deterministic system of non-interacting agents with no Brownian motion (and again we switch notations of time-dependence in deterministic quantities, $\mathbf{x}^{j}(t),m^{j}(t),\overline{F}^{N}(t)$ etc.)
@@
The essential role of communication was already highlighted in the context of SBGD method, [21]: once we decouple the step-size and mass, setting a uniform step-size in (SBGD), that is, $\alpha(m)\equiv 1$, then the success rate of the resulting non-communicating swarm-based method decreases dramatically. The system (4.1) provides yet another realization of the essential role of communication. We note that since all particles are “descending,” it is also straightforward to show that the collective mean objective $\overline{F}^{N}(t)$ is always decreasing: $\frac{\textnormal{d}}{\textnormal{d}t}\overline{F}^{N}(t)\leqslant 0$. See Section A.1 for the proof. However, even though $\overline{F}^{N}(t)$ continues to decay in time, it does not guarantee the desired convergence to the global minimum, $\liminf_{t\rightarrow\infty}\overline{F}^{N}(t)=\min_{\mathbf{x}}F(\mathbf{x})$. This is due to the passive nature of the gradient descent of the agents. In particular, it is likely that all agents will become trapped at local minima or saddle points, especially when the number of particles is not sufficiently large. Turning on the communication term that adjusts strength of Brownian motion, as done in (SSA), allows particles situated at local minima to experience increased randomness. This increased randomness facilitates their escape from local basins. With the communication and the randomness turned off, we have the following:
@@ 5.1. Convergence to the mean-field limit
Theorem 2.1 states that the empirical distribution $\mu^{N}_{t}$ that assembles all samples that follow the coupled SDE system (SSA), is close to the limiting distribution $\mu$ that solves the PDE (2.4a).
There are a few different but related approaches for justifying mean-field limits, including the coupling method, the energy/entropy estimates, the utilization of BBGKY (Bogoliubov–Born–Green–Kirkwood–Yvon hierarchy) hierarchies and running the tightness argument to confine a sequence of measures for the convergence; see references [22, 28]. Here we employ the coupling method, introduced in [28]. The coupling method places a stronger requirement on the field, but the machinery is rather simple to use, and it has the advantage of providing precise convergence rate in terms of the number of particles, $N$.
@@
(5.2) $\widetilde{\mu}^{N}_{t}=\frac{1}{N}\sum_{i}\delta_{\widetilde{\mathbf{x}}^{j}_{t}}(\mathbf{x})\otimes\delta_{\widetilde{m}^{j}_{t}}(m)\,,\quad\text{and}\quad\widetilde{F}^{N}_{t}=\frac{\frac{1}{N}\sum_{j}\widetilde{m}^{j}_{t}F\left(\widetilde{\mathbf{x}}^{j}_{t}\right)}{\mathbb{E}_{\mu}[m]}\,.$
We note that this system has a clear similarity with the original SDE (SSA) but is passive, in the sense that the dynamics of $\widetilde{m}^{j}_{t}$ is driven by ¯$\overline{\mathscr{F}}^{\mu}$ instead of $\widetilde{F}^{N}_{t}$. While the dynamics of $\widetilde{\mathbf{x}}_{t}$ agrees with that of $\mathbf{x}_{t}$, the mass changes call for different rates. In particular, in (SSA), we compare $F(\mathbf{x}_{t})$ with $\overline{F}^{N}_{t}$ that is self-generated in (2.2), and here, $F(\widetilde{\mathbf{x}}_{t})$ is compared with ¯$\overline{\mathscr{F}}^{\mu}$ defined from the underlying PDE system (2.4a). As a consequence, this new $\{\widetilde{\mathbf{x}}^{j}_{0}=\mathbf{x}^{j}_{0}\,,\widetilde{m}^{j}_{0}=m^{j}_{0}\}$ system gets passively pushed forward by the underlying $\mu$.
@@
    $W_{2}(\mu_{t},\widetilde{\mu}^{N}_{t})\to 0\,,\quad\text{in probability as }N\to\infty\,.$
@@
Suppose Assumption 2.1 holds. Let $\{\widetilde{\mathbf{x}}^{j}_{t}\,,\widetilde{m}^{j}_{t}\}_{j=1}^{N}$ be the system defined in (5.1), and let $\{{\mathbf{x}}^{j}_{t}\,,{m}^{j}_{t}\}_{j=1}^{N}$ solve (SSA). Suppose the two systems have the same initial conditions, then
@@
    (5.3) $W_{2}(\mu^{N}_{t},\widetilde{\mu}^{N}_{t})\to 0\,,\quad\text{in probability as}N\to\infty\,.$
@@
    (5.4) $\mathbb{E}(|\widetilde{F}^{N}_{t}-\overline{F}^{N}_{t}|)<\frac{C}{\sqrt{N}}\,.$
@@
With these two lemmas, Theorem 2.1 is a straightforward corollary, calling triangle inequalities:
$W_{2}(\mu_{t},\mu^{N}_{t})\leqslant W_{2}(\mu_{t},\widetilde{\mu}^{N}_{t})+W_{2}(\widetilde{\mu}^{N}_{t},\mu^{N}_{t})\,,$
and
$\mathbb{E}(|\overline{\mathscr{F}}^{\mu}_{t}-\overline{F}^{N}_{t}|)\leqslant\mathbb{E}(|\overline{\mathscr{F}}^{\mu}_{t}-\widetilde{F}^{N}_{t}|)+\mathbb{E}(|\widetilde{F}^{N}_{t}-\overline{F}^{N}_{t}|)\,.$
Here we omit the details and proceed to provide the proof of the two lemmas.
Proof of Lemma 5.1.
Calling Theorem 1 in [9],
@@
with $\theta=\max\{1/2,1/d\}$. This implies
$W_{2}(\mu_{t},\widetilde{\mu}^{N}_{t})\to 0\,,\quad\text{in probability as }N\to\infty\,.$
@@
The proof for Lemma 5.2 is the standard coupling method. This amounts to comparing the original SDE (SSA) with the auxiliary SDE (5.1). In particular, we note that the $\mathbf{x}$ component of the two systems are almost identical:
@@
Proof of Lemma 5.2.
Noting that $\mu^{N}_{t}=\frac{1}{N}\sum_{j}\delta_{\mathbf{x}^{j}_{t}}(\mathbf{x})\otimes\delta_{m^{j}_{t}}(m)$ and $\widetilde{\mu}^{N}_{t}=\frac{1}{N}\sum_{j}\delta_{\widetilde{\mathbf{x}}^{j}_{t}}(\mathbf{x})\otimes\delta_{\widetilde{m}^{j}_{t}}(m)$, then recall the definition of the Wasserstein metric, we have:
@@
Define:
$\Omega=\left\{\left|\frac{1}{N}\sum_{j=1}^{N}m^{j}_{0}-\mathbb{E}_{\mu_{0}}(m)\right|\leqslant\frac{1}{2}\right\}\,,$
With Chebyshev’s inequality and boundedness of $m$ from Assumption 2.1, we obtain
$\mathbb{P}\left(\Omega^{c}\right)\leqslant\frac{4\mathbb{E}_{\mu_{0}}(m^{2})}{N}\leqslant\frac{C}{N}\,,\quad\text{and thus accordingly}\quad\mathbb{P}\left(\Omega\right)\geqslant 1-\frac{C}{N}\geqslant\frac{1}{2}\,.$
Define
(5.6) $G(t)=\mathbb{E}\left(\frac{1}{N}\sum^{N}_{j=1}\left[|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}+\left|m^{j}_{t}-\widetilde{m}^{j}_{t}\right|^{2}\right]\middle|\Omega\right)\,.$
If we can show $G(t)\sim 1/N$, then conditioned on $\Omega$, a set whose probability goes to $1$, $W_{2}(\widetilde{\mu}^{N}_{t},\mu^{N}_{t})^{2}\to 0$ as $N\to\infty$, proving (5.3).
Similarly, to estimate (5.4), we note:
(5.7) $|\widetilde{F}^{N}_{t}-\overline{F}^{N}_{t}|$
$\leqslant$ $\left|\frac{\frac{1}{N}\sum^{N}_{j=1}\left(\widetilde{m}^{j}_{t}F\left(\widetilde{\mathbf{x}}^{j}_{t}\right)-m^{j}_{t}F(\mathbf{x}^{j}_{t})\right)}{\mathbb{E}_{\mu}[m]}\right|+\left|\left(\frac{1}{N}\sum^{N}_{j=1}m^{j}_{t}F(\mathbf{x}^{j}_{t})\right)\left(\frac{1}{\mathbb{E}_{\mu}[m]}-\frac{1}{\frac{1}{N}\sum m^{j}_{t}}\right)\right|$
$\underset{\rm(i)}{\leqslant}$ $C\left[\left|\frac{1}{N}\sum^{N}_{j=1}\left(m^{j}_{t}-\widetilde{m}^{j}_{t}\right)F(\widetilde{\mathbf{x}}^{j}_{t})+\frac{1}{N}\sum^{N}_{j=1}m^{j}_{t}\left(F(\mathbf{x}^{j}_{t})-F\left(\widetilde{\mathbf{x}}^{j}_{t}\right)\right)\right|+\left|\frac{\frac{1}{N}\sum m^{j}_{t}-\mathbb{E}_{\mu}[m]}{\mathbb{E}_{\mu}[m]\left(\frac{1}{N}\sum m^{j}_{t}\right)}\right|\right]$
=$\underset{\rm(ii)}{=}$ $C\left[\left|\frac{1}{N}\sum^{N}_{j=1}\left(m^{j}_{t}-\widetilde{m}^{j}_{t}\right)F(\widetilde{\mathbf{x}}^{j}_{t})+\frac{1}{N}\sum^{N}_{j=1}m^{j}_{t}\left(F(\mathbf{x}^{j}_{t})-F\left(\widetilde{\mathbf{x}}^{j}_{t}\right)\right)\right|+\left|\frac{\frac{1}{N}\sum m^{j}_{0}-\mathbb{E}_{\mu_{0}}[m]}{\mathbb{E}_{\mu_{0}}[m]\left(\frac{1}{N}\sum m^{j}_{0}\right)}\right|\right]$
$\underset{\rm(iii)}{\leqslant}$ $C\left[\sqrt{\frac{1}{N}\sum^{N}_{j=1}\left[|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}+\left|m^{j}_{t}-\widetilde{m}^{j}_{t}\right|^{2}\right]}+\left|{\frac{1}{N}\sum m^{j}_{0}-\mathbb{E}_{\mu_{0}}[m]}\right|\right]\,,$
Here in (i) we used the upper boundedness of both $m$ [^5] Firstly we notice that, according to the Boundedness in $m$ condition in Assumption 2.1, the initial data for $m^{j}_{t}$ is bounded by $M_{\max}$, meaning $|m^{j}_{0}|<M_{\max}$ for all $j$ at the initial time. Given that $|F|\leqslant D$, then according to the $m^{j}_{t}$ equation in (SSA), we call Grönwall inequality to obtain that (5.8) $|m^{j}_{t}|\leqslant M_{\max}e^{2Dt}$ for all time. The same argument applies to bound $\widetilde{m}^{j}_{t}$. We denote the upper bound $c$ that depends on $M_{\max}$, $D$ and $t$. and $F$. In (ii), we apply the conservation of mass: $\sum m^{j}_{t}=\sum m^{j}_{0}$ and $\mathbb{E}_{\mu}[m]=\mathbb{E}_{\mu_{0}}[m]$. In (iii), we used the triangle inequality, Cauchy-Schwarz inequality, and Lipschitz continuity of $F$. Consequently, the constant $C$ depends on $D$, $L$, and $M_{\max}$ (from Assumption 2.1).
To insert it to (5.4), then:
$\mathbb{E}(|\widetilde{F}^{N}_{t}-\overline{F}^{N}_{t}|)\leqslant\mathbb{E}\left(|\widetilde{F}^{N}_{t}-\overline{F}^{N}_{t}|\big{|}\Omega\right)\mathbb{P}(\Omega)+\mathbb{E}\left(|\widetilde{F}^{N}_{t}-\overline{F}^{N}_{t}|\big{|}\Omega^{c}\right)\mathbb{P}(\Omega^{c})\,.$
So combining with (5.7), we rewrite (5.4):